OpenStudy (anonymous):

can t squared plus 5t be factored?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yup. Since both terms have t, you can take out one t , giving you t(t+5)

OpenStudy (unklerhaukus):

\[t^2+5t=t\times t+t\times5=t(t+5)\]